# Insert In Style: A Zero-Shot Generative Framework for Harmonious Cross-Domain Object Composition

### v1

@@ Abstract
Reference-based object composition methods fail when inserting real-world objects into stylized domains. This under-explored problem is currently split between practical “blenders” that lack generative fidelity and “generators” that require impractical, per-subject online finetuning. In this work, we introduce Insert In Style, the first zero-shot generative framework that is both practical and high-fidelity. Our core contribution is a unified framework with two key innovations: (i) a novel multi-stage training protocol that disentangles representations for identity, style, and composition, and (ii) a specialized masked-attention architecture that surgically enforces this disentanglement during generation. This approach prevents the concept interference common in general-purpose, unified-attention models. Our framework is trained on a new 100100k sample dataset, curated from a novel data pipeline. This pipeline couples large-scale generation with a rigorous, two-stage filtering process to ensure both high-fidelity semantic identity and style coherence. Unlike prior work, our model is truly zero-shot and requires no text prompts. We also introduce a new public benchmark for stylized composition. We demonstrate state-of-the-art performance, significantly outperforming existing methods on both identity and style metrics, a result strongly corroborated by user studies.
@@ 1 Introduction
Reference-based object composition, focused on the task of inserting a specific object into a scene [8, 22, 47], is a fundamental challenge in computer vision. Recent methods like DreamFuse [15], AnyDoor [6] and IMPRINT [43] have achieved remarkable realism. However, these models are trained almost exclusively on photorealistic data and fail spectacularly when composing objects into stylized domains like paintings, sketches, or digital art—a vast and common use case.
@@
In this work, we introduce Insert In Style, the first framework to solve this challenge. Our core methodological contribution is two-fold: a novel training protocol and a specialized attention architecture. First, we propose a three-stage training protocol to explicitly disentangle representations: (a) a reference object encoder to learn robust identity, (b) a spatial style encoder to learn generalizable style, and (c) a final composition stage. Second, to surgically enforce this disentanglement at inference time and prevent feature-bleed, we introduce a novel masked-attention mechanism. This specialized architecture stands in direct contrast to general-purpose, unified-attention models and is key to balancing our competing objectives.
@@
To power this framework, we introduce a 100100k sample training corpus, created via a novel data pipeline that couples large-scale, multi-method generation with a robust, two-stage filtering process. We objectively calibrate our filtering thresholds on a $1,000$ sample, human-annotated validation set, which ensures our final dataset meets a high standard for both identity preservation and style coherence.
@@
    A novel generative framework featuring: (i) a three-stage training protocol that learns disentangled encoders for identity and style, and (ii) a masked-attention architecture that prevents feature-bleed between these competing conditions during composition.
@@
    The largest-scale dataset for this task (100100k samples), curated by a novel, two-stage pipeline that is rigorously calibrated on human annotations to ensure both semantic identity and style coherence. We will make both the dataset and protocol public.
@@ 2.3 Style Transfer
Figure 3: Dataset Pipeline. (a) Generation: We create a large-scale, diverse raw corpus by applying a mix of state-of-the-art stylization methods (FLUX.1-Kontext [20], CSGO [51], and CAST [54]). (b) Filtering: Our raw dataset is then refined by our rigorous two-stage filtering process. The Identity Consistency filter prunes samples with semantic drift in the subject region, while the Style Coherence filter removes aesthetic mismatches between the subject region and its surrounding background, together ensuring a high-fidelity dataset.
@@ 3 Dataset Generation
Our framework is powered by a new, large-scale dataset. This section details our data curation methodology, which consists of a large-scale corpus generation and our novel filtering pipeline. The entire process is illustrated in Fig. 3. This data-centric approach is the foundation for our model’s zero-shot generative capabilities.
@@ 3.1 Data Generation Pipeline
We build upon the DreamFuse dataset [15], which provides high-quality $\{I_{f},I_{c},I_{m}\}$ triplets. Upon inspection, we observed that a subset of these triplets contains a semantic mismatch between the reference $I_{f}$ and the object in the composite image $I_{c}$. To ensure the quality of our base data, we first proactively filter out these mismatched samples using CLIP-based similarity [12] between the foreground reference $I_{f}$ and the masked object region in $I_{c}$. From this cleaned set, we then select foreground references from relevant classes (e.g., object, handheld, animal, pet, and product). This curation process yields our final base dataset, $D_{b}$, of approximately $40,000$ high-quality triplets.
@@
To generate a training corpus with maximum fidelity and style diversity, we employed a principled, multi-pronged strategy. Our primary generative pipeline is built on FLUX.1-Kontext [20], which we chose for its state-of-the-art performance in structure-preserving stylization [21], [35]. We trained $20$ custom LoRA [14] modules for it on high-quality external style datasets [4] to create variants for popular and complex archetypes, such as Ghibli, Watercolor, and Chinese Ink.
To broaden the style diversity beyond these $20$ archetypes and ensure our model generalizes, we supplemented this pipeline with state-of-the-art reference-based methods. We employed CSGO [51] and CAST [54], which our empirical analysis showed are highly effective at preserving the subject’s spatial integrity and visual identity. We paired these methods with the Style3030k collection [24], a diverse benchmark covering $1,120$ fine-grained style categories.
This combined generation process yields an initial corpus $D_{b}^{sty}$ of approximately 150150k stylized compositions. Each sample in this final dataset consists of a $\{I_{f},I_{c},I_{m},I_{s}\}$ quadruplet, where $I_{s}$ is the stylized composite image.
@@ 3.2 Filtering Methods
This large, raw dataset inevitably contains a spectrum of failure cases, including, subject’s semantic identity drift, and local style incoherence. A rigorous filtering stage is therefore essential to curate a high-quality corpus. We propose a novel, two-stage hybrid filtering pipeline to ensure both semantic identity and style coherence.
Filter 1: Identity Consistency Filtering.
After obtaining the raw stylized dataset $D_{b}^{sty}$, it is important to ensure that the identity of the subject remains consistent between the composite image $I_{c}$ and the stylized composite image $I_{s}$. For this, we utilize the CLIP [12] score, which measures semantic similarity, and DINO [30], which measures structural similarity. First, we define an operation $\mathcal{C}(I,I_{m})$ which, for a given image $I$ and mask $I_{m}$, crops the region of $I$ indicated by $I_{m}$. For every sample of $D_{b}^{sty}$, we calculate $S_{\text{clip}}$ and $S_{\text{dino}}$ as follows:
$S_{\text{clip}}$ $=\text{CLIPSim}(\mathcal{C}(I_{s},I_{m}),\mathcal{C}(I_{c},I_{m}))$ (1)
$S_{\text{dino}}$ $=\text{DINO}(\mathcal{C}(I_{s},I_{m}),\mathcal{C}(I_{c},I_{m}))$ (2)
We use a held-out validation set of size $1,000$ from $D_{b}^{sty}$. We manually classify the pairs $I_{c}$, $I_{s}$ based on whether the object identities match. For these same pairs, we calculate the $S_{\text{clip}}$ and $S_{\text{dino}}$ scores. We perform a grid search over the range of the respective scores to find thresholds $\mathcal{T}_{\text{clip}}$ and $\mathcal{T}_{\text{dino}}$ that maximize precision while keeping the rejection rate below a specific limit. The details of this process are presented in the supplementary material. These thresholds are used to filter $D_{b}^{sty}$; we accept only those samples that satisfy $S_{\text{clip}}>\mathcal{T}_{\text{clip}}$ and $S_{\text{dino}}>\mathcal{T}_{\text{dino}}$.
Figure 4: Qualitative samples from our Insert In Style Dataset. Spanning 100100k samples and $1,140$ unique styles, it is the largest-scale corpus for this task. Each <<Subject, Composite, Stylized Composite>> triplet provides the strong, aligned supervision required to train robust, cross-domain insertion models.
@@
Dataset Venue Task Type Foreground Ref. Composite Image Foreground Ref. Mask Stylized Composite Image # Styles # Samples Style-30K [24] ECCV $2024$ Stylization $1120$ 3030k Wiki-Art [39] arXiv $2015$ Stylization $27$ 5757k ArtBench [25] arXiv $2022$ Stylization $10$ 6060k OmniConsistency [44] NeurIPS $2025$ Stylization $22$ $2600$ OmniStyle [50] CVPR $2025$ Ref. based Stylization (has Style Ref.) $1000$ 150150k DreamFuse [15] ICCV $2025$ Object Insertion - 8484k Insert In Style - Cross-domain Object Insertion $\mathbf{1140}$ $\mathbf{100}$k
Filter 2: Style Coherence Filtering.
While identity-based filtering ensures semantic consistency, it does not guarantee stylistic coherence between the stylized object and its background. We observe that occasionally, stylization disproportionately affects some regions including the subject region, resulting in a perceptually inconsistent stylization between the subject and the rest of the image $I_{s}$. To ensure visual harmony, we use CSD [41], which measures the similarity of style characteristics between two images. We use CSD to calculate style similarity between the subject region and the remaining areas of the image $I_{s}$. For every sample $\{I_{s},I_{m}\}$ from $D_{b}^{sty}$, we calculate the CSD score:
$S_{\text{csd}}=\text{CSD}(\mathcal{C}(I_{s},I_{m}),\mathcal{R}(\mathcal{C}(I_{s},1-I_{m}))),$ (3)
where $\mathcal{R}$ is an operation that copies patches of size $64\times 64$ from retained regions to masked-out regions.The threshold $\mathcal{T}_{\text{csd}}$ for filtering samples is determined in a similar way as described in the identity filtering stage. Using this threshold, we only accept those samples from $D_{b}^{sty}$ that satisfy $S_{\text{csd}}>\mathcal{T}_{\text{csd}}$.
This two-stage filtering process reduces our 150150k generated samples to the final 100100k high-quality, identity-preserving, and style-coherent training corpus. We provide qualitative samples in Fig. 4 and a detailed comparison to existing datasets in Tab. 1 that establishes our corpus as the first and largest to provide the aligned $\{I_{f},I_{s},I_{m}\}$ triplets essential for this task.
@@ 4 Proposed Method
The core challenge in cross-domain object composition is to balance two competing objectives: identity preservation (which requires preserving the object’s features) and style harmonization (which requires transforming them). This presents a unique challenge for general-purpose, unified-attention models [45], which process all conditional signals jointly. This joint processing, while powerful, is not explicitly designed to manage competing signals, creating a potential risk of concept interference or feature bleed, where the strong style signal may corrupt identity features, or vice-versa. To solve this, we introduce a novel framework that is a two-fold contribution:
@@
    A three-stage training protocol that explicitly disentangles these competing concepts by pre-training specialized, independent encoders for identity and style.
@@
    A specialized masked-attention architecture that surgically enforces this disentanglement during the final composition stage, preventing concept interference.
@@ 4.1 Model Architecture
Our architecture extends the FLUX.11-dev [19] framework, a dual-branch DiT [32]. It operates on latent representations: a VAE [17] encodes images into latents $Z_{0}\in\mathbb{R}^{H\times W\times C}$, and a T55 encoder [34] produces text embeddings $Z_{c}\in\mathbb{R}^{L\times D}$. The model is trained as a rectified flow [26] to predict the velocity vector $v_{\theta}$ of a flow matching a linear interpolation between noise $Z_{1}\sim\mathcal{N}(0,I)$ and the target image latent $Z_{0}$. We define the path as $Z_{t}=t\cdot Z_{0}+(1-t)\cdot Z_{1}$, where $t\in[0,1]$. The target velocity is $v^{*}=Z_{0}-Z_{1}$, and the objective is the $L_{2}$ loss:
@@
To handle our competing conditions, we extend the FLUX.11-dev [19] architecture with two additional, parameter-efficient conditioning branches. The full model processes four parallel token sequences:
@@
These are then concatenated with the image ($Q_{t},K_{t},V_{t}$) and text ($Q_{c},K_{c},V_{c}$) tokens for the shared self-attention operation [45], which we detail in Stage $3$.
4.2 Multi-stage Training Protocol
@@
We train the LoRA adapters of our architecture in three distinct stages to incrementally build and align the required representations. Our protocol uses the same frozen, pre-trained FLUX.11-dev [19] model weights as the foundation for three distinct training stages. We first train two specialist encoders independently and in parallel (Stages $1$, 22), then assemble them to train a final compose model (Stage $3$).
Stage 1: Reference Object Encoder.
@@
We first attach a new Identity Branch ($Z_{ref}$) to the frozen, pre-trained FLUX.11-dev [19] base model and train only its LoRA adapters. The model is trained on the Subjects200200K Collection $2$ dataset [45], which contains paired reference objects and their corresponding composed scenes. We use text prompts that implicitly reference the subject (e.g., “a photo of this item on a table”), forcing the model to ground the subject’s identity in the visual features of $Z_{ref}$. The model is trained to reconstruct the full scene using $\mathcal{L}_{\text{flow}}$ (Eq. 4), conditioned on $Z_{t}$, $Z_{ref}$, and $Z_{c}$.
Stage 2: Spatial Style Encoder.
@@
Independently, we attach a new Style Branch ($Z_{style}$) to the same frozen, pre-trained FLUX.11-dev [19] base model and train only its LoRA adapters. This parallel process ensures the style representations are learned completely independently from the identity representations. The model is trained on a diverse corpus of $70,000$ images, including $40,000$ stylized scenes from OmniStyle [50], $15,000$ from StyleBooth [11], and $15,000$ real-world images. We train the model on a style-aware inpainting task. Given an image latent $Z_{i}$ and a binary mask $M$, we define the style context as the unmasked tokens $Z_{\text{style}}=Z_{i}\odot(1-M)$ and the noisy target tokens as $Z_{t}$ (the masked region $Z_{i}\odot M$, noised). The model is trained to denoise $Z_{t}$ conditioned on $Z_{\text{style}}$ and a text prompt $Z_{c}$.
Stage 3: Composition with Masked Attention.
Finally, we assemble the complete compositional model. We load the frozen, pre-trained FLUX.11-dev [19] base model and attach both the pre-trained and frozen Identity Branch from Stage $1$ and the pre-trained and frozen Style Branch from Stage $2$. We now introduce and train a new set of LoRA adapters, $\Delta W_{\theta}$, on the main branch ($Z_{t}$). This stage is trained on our new Insert In Style Dataset containing 100100k samples (see Sec. 3).

### v2

@@ Abstract
Reference-based object composition involves integrating foreground reference image with background scene to produce harmonious fused image. This task becomes particularly challenging in cross-domain scenarios, where models must balance preserving the reference object’s identity while harmonizing them to match stylized environments. This under-explored problem is currently split between practical “blenders” that lack generative fidelity and “generators” that require impractical, per-subject online finetuning. In this work, we introduce Insert In Style, the first zero-shot generative framework that is both practical and high-fidelity. Our core contribution is a unified framework with two key innovations: (i) a novel multi-stage training protocol that disentangles representations for identity, style, and composition, and (ii) a specialized masked-attention architecture that surgically enforces this disentanglement during generation (iii) A prior preservation objective that keeps learned identity and style priors intact. By design, this approach mitigates concept interference typical in unified-attention architectures while ensuring robust generalization across diverse references and styles. Our framework is trained on a new 115115k sample dataset, curated from a novel data pipeline. This pipeline couples large-scale generation with a rigorous, iterative human-in-the-loop filtering process to ensure both high-fidelity semantic identity and style coherence. Unlike prior work, our model is truly zero-shot and requires no text prompts. We also introduce a new public benchmark for stylized composition. We demonstrate state-of-the-art performance, significantly outperforming existing methods on both identity and style metrics, a result strongly corroborated by user studies.
@@ 1 Introduction
Reference-based object composition, focused on the task of inserting a specific object, identified by the reference image, in a scene at a specific location, identified by provided user mask, [8, 22, 47], is a fundamental challenge in computer vision. Recent methods like DreamFuse [15], AnyDoor [6] and IMPRINT [43] have achieved remarkable realism. However, these models are trained almost exclusively on photorealistic data and fail spectacularly when composing objects into stylized domains like paintings, sketches, or digital art—a vast and common use case.
@@
In this work, we introduce Insert In Style, the first framework to solve this challenge. Our core methodological contribution is two-fold: First, we propose a multi-stage training protocol to explicitly disentangle representations: (a) a reference object encoder to learn robust identity, (b) a spatial style encoder to learn generalizable style, and (c) a final composition stage. Second, A novel approach to compose the trained competing conditions with attention disentanglement and prior preservation losses. This specialized architecture stands in direct contrast to general-purpose, unified-attention models and is key to balancing our competing objectives.
@@
To power this framework, we introduce a 115115k sample training corpus, created via a novel data pipeline that couples large-scale, multi-method generation with a robust, filtering process. Our iterative human-in-the-loop data filtering process involves ensemble of identity and style evaluation functions, which ensures our final dataset meets a high standard for both identity preservation and style coherence.
@@
    A novel generative framework featuring: (i) a three-stage training protocol that learns disentangled encoders for identity and style, and (ii) a masked-attention architecture that prevents feature-bleed between these competing conditions during composition (iii) a prior preservation objective that keeps learned identity and style priors intact.
@@
    The largest-scale dataset for this task (115115k samples), curated by a novel, two-stage pipeline that is rigorously calibrated on human annotations to ensure both semantic identity and style coherence. We will make both the dataset and protocol public.
@@ 2.3 Style Transfer
Figure 3: Dataset Pipeline. (a) Generation: We create a large-scale, diverse raw corpus by applying a mix of state-of-the-art stylization methods (FLUX.1-Kontext [20], CSGO [51], and CAST [54]). (b) Filtering: Our raw dataset is then refined by our rigorous iterative filtering process involving identity consistency evaluation and style coherence evaluation with human-in-the-loop (see Sec. 3.2).
@@ 3 Dataset Generation
Traditional dataset curation approaches for reference-based object composition are limited to in-domain composition [6, 15, 5]. To enable cross-domain compositional generalization, we propose a novel scalable approach built upon existing in-domain composition data. This section details our data curation methodology, which consists of a large-scale corpus generation and our novel filtering pipeline. The entire process is illustrated in Fig. 3. This data-centric approach is the foundation for our model’s zero-shot generative capabilities.
@@ 3.1 Data Generation Pipeline
We build upon the DreamFuse dataset [15], which curated diverse $\{I_{f},I_{c},I_{m}\}$ triplets through reference based generation. Upon inspection, we observed that a subset of these triplets contains a semantic mismatch between the reference $I_{f}$ and the object in the composite image $I_{c}$. To ensure the quality of our base data, we first proactively filter out these mismatched samples using CLIP-based similarity [12] between the foreground reference $I_{f}$ and the masked object region in $I_{c}$. From this cleaned set, we then select foreground references from relevant classes (e.g., object, handheld, animal, pet, and product). This curation process yields our final base dataset, $\mathcal{D}^{base}$, of approximately $40,000$ high-quality triplets. Our core strategy for extending in-domain data to cross-domain involves performing diverse identity-preserving stylisation on in-domain composite images $I_{c}$ to generate its stylised variant $I_{s}$.
@@
Our primary generative pipeline is built on FLUX.1-Kontext [20], which we chose for its state-of-the-art performance in structure-preserving stylization [21], [35]. We trained $20$ custom LoRA [14] modules for it on high-quality external style datasets [4] to create variants for popular and complex archetypes, such as Ghibli, Watercolor, and Chinese Ink.
To broaden the style diversity beyond these $20$ archetypes and ensure our model generalizes, we supplemented this pipeline with state-of-the-art reference-based methods. We employed CSGO [51] and CAST [54], which our empirical analysis showed are highly effective at preserving the subject’s structural integrity and visual identity. We paired these methods with the Style3030k collection [24], a diverse benchmark covering $1,120$ fine-grained style categories.
This combined generation process yields an initial corpus $\mathcal{D}^{sty}$ of approximately 350350k stylized compositions. Each sample in this final dataset consists of a $\{I_{f},I_{c},I_{m},I_{s}\}$ quadruplet, where $I_{s}$ is the stylized composite image.
@@ 3.2 Filtering Methods
This large, raw dataset inevitably contains a spectrum of failure cases, including, subject’s semantic identity drift, and local style incoherence. A rigorous filtering stage is therefore essential to curate a high-quality corpus. We propose a novel, hybrid filtering pipeline to ensure both identity consistency and style coherence.
Identity Consistency Evaluation.
After obtaining the raw stylized dataset $\mathcal{D}^{sty}$, it is important to ensure that the identity of the subject remains consistent between the composite image $I_{c}$ and the stylized composite image $I_{s}$. For this, we utilize the CLIP [12] score, which measures semantic similarity, and DINO [30], which measures structural similarity. First, we define an operation $\mathcal{C}(I,I_{m})$ which, for a given image $I$ and mask $I_{m}$, crops the region of $I$ indicated by $I_{m}$. For every sample of $\mathcal{D}^{sty}$, we calculate $\phi_{\text{clip}}$ and $\phi_{\text{dino}}$ as follows:
$\phi_{\text{clip}}$ $=\text{CLIPSim}(\mathcal{C}(I_{s},I_{m}),\mathcal{C}(I_{c},I_{m}))$ (1)
$\phi_{\text{dino}}$ $=\text{DINO}(\mathcal{C}(I_{s},I_{m}),\mathcal{C}(I_{c},I_{m}))$ (2)
@@
Dataset Venue Task Type Foreground Ref. Composite Image Foreground Ref. Mask Stylized Composite Image # Styles # Samples Style-30K [24] ECCV $2024$ Stylization $1120$ 3030k Wiki-Art [39] arXiv $2015$ Stylization $27$ 5757k ArtBench [25] arXiv $2022$ Stylization $10$ 6060k OmniConsistency [44] NeurIPS $2025$ Stylization $22$ $2600$ OmniStyle [50] CVPR $2025$ Ref. based Stylization (has Style Ref.) $1000$ 150150k DreamFuse [15] ICCV $2025$ Object Insertion - 8484k Insert In Style - Cross-domain Object Insertion $\mathbf{1140}$ $\mathbf{115}$k
Style Coherence Evaluation
While identity-based evaluation ensures semantic consistency, it does not guarantee stylistic coherence between the stylized object and its background. We observe that occasionally, stylization disproportionately affects some regions including the subject region, resulting in a perceptually inconsistent stylization between the subject and the rest of the image $I_{s}$. To ensure visual harmony, we use CSD [41], which measures the similarity of style characteristics between two images. We use CSD to calculate style similarity between the subject region and the remaining areas of the image $I_{s}$. For every sample $\{I_{s},I_{m}\}$ from $\mathcal{D}^{sty}$, we calculate the CSD score:
$\phi_{\text{csd}}=\text{CSD}(\mathcal{C}(I_{s},I_{m}),\mathcal{P}(\mathcal{C}(I_{s},1-I_{m}))),$ (3)
where $\mathcal{P}$ is an operation that copies patches of size $64\times 64$ from retained regions to masked-out regions.
Iterative Human-in-the-loop Multidimensional Data Filtering
To ensure the high alignment of our dataset across identity and style, we implement a progressive multidimensional filtering strategy. We define an ensemble of scoring functions $\Phi=\{\phi_{\text{CLIP}},\phi_{\text{DINO}},\phi_{\text{CSD}}\}$ to evaluate the semantic, structural, and stylistic integrity of each instance $x\in\mathcal{D}^{sty}$.
The refinement process is formulated as an iterative distillation. At each iteration $t$, we identify a rejection set $\mathcal{R}_{t}$ comprised of the union of the lowest deciles across all scoring dimensions:
$\mathcal{R}_{t}=\bigcup_{\phi\in\Phi}\{x\in\mathcal{D}_{t}\mid\phi(x)\leq Q_{0.1}(\phi)\}$ (4)
where $Q_{0.1}(\phi)$ represents the $10^{th}$ percentile of the score distribution for metric $\phi$ within the current subset $\mathcal{D}_{t}^{sty}$. The dataset is subsequently filtered via $\mathcal{D}_{t+1}^{sty}=\mathcal{D}_{t}^{sty}\setminus\mathcal{R}_{t}$.
Rather than employing a fixed iteration count, we adopt a human-centric convergence heuristic to balance dataset purity with sample diversity. After each pruning step, the rejection set $\mathcal{R}_{t}$ undergoes expert qualitative assessment. The process terminates at iteration $T$ when the samples in $\mathcal{R}_{T}$ exhibit high perceptual quality, specifically regarding identity preservation and stylistic fidelity. This filtering process reduces generated samples to the final 115115k high-quality, identity-preserving, and style-coherent training corpus. We provide more details on the dataset generation, filtering process and training data composition in the supplementary. Tab. 1 shows a detailed comparison to existing datasets establishing our corpus as the first and largest to provide the aligned $\{I_{f},I_{c},I_{m},I_{s}\}$ quadruplets essential for this task.
@@ 4 Proposed Method
The core challenge in cross-domain object composition is to balance two competing objectives: identity preservation (which requires preserving the object’s features) and style harmonization (which requires transforming them). This presents a unique challenge for general purpose multi condition adapter fusion methods [45], where all independently trained conditions are trained with same generative objective. This joint processing, while powerful, is not explicitly designed to manage competing signals, creating a potential risk of concept interference or feature bleed, where the strong style signal may corrupt identity features, or vice-versa. To solve this, we introduce a novel framework that is a two-fold contribution:
@@
    A multi stage training protocol that explicitly disentangles these competing concepts by pre-training specialized, independent encoders for identity and style objectives.
@@
    A novel approach to compose the pre-trained competing conditions with attention disentanglement and prior preservation.
@@ 4.1 Model Architecture
Our proposed framework built atop pre-trained DiTs, using FLUX.11-dev as our primary base model [19]. Notably, our formulation is backbone agnostic and can be extended to modern T2I or instruction-based editing DiTs [21, 32]. FLUX.11-dev operates on latent representations: a VAE [17] encodes images into latents $Z_{0}\in\mathbb{R}^{H\times W\times C}$, and a T55 encoder [34] produces text embeddings $Z_{c}\in\mathbb{R}^{L\times D}$. The model is trained as a rectified flow [26] to predict the velocity vector $v_{\theta}$ of a flow matching a linear interpolation between noise $Z_{1}\sim\mathcal{N}(0,I)$ and the target image latent $Z_{0}$. We define the path as $Z_{t}=t\cdot Z_{0}+(1-t)\cdot Z_{1}$, where $t\in[0,1]$. The target velocity is $v^{*}=Z_{0}-Z_{1}$, and the flow matching objective:
@@
To handle our competing conditions, we extend the FLUX.11-dev [19] architecture with two additional, parameter-efficient conditioning branches [45] . The full model processes four parallel token sequences:
@@
These are then concatenated with the image ($Q_{t},K_{t},V_{t}$) and text ($Q_{c},K_{c},V_{c}$) tokens for the shared self-attention operation [45].
4.2 Conditional Branch Training
@@
We train the LoRA adapters of our architecture to incrementally build and align the required representations. Our protocol uses the same frozen, pre-trained FLUX.11-dev [19] model weights as the foundation for three distinct training stages. We first train two specialist encoders independently and in parallel, then assemble them to train a final compose model Insert In Style.
Reference Object Encoder.
@@
We first attach a new Identity Branch ($Z_{ref}$) to the frozen, pre-trained FLUX.11-dev [19] base model and train only its LoRA adapters ($\Delta W_{\text{ref}}$). The model is trained on the Subjects200200K Collection $2$ dataset [45], which contains paired reference objects and their corresponding composed scenes. We use text prompts that implicitly reference the subject (e.g., “a photo of this item on a table”), forcing the model to ground the subject’s identity in the visual features of $Z_{ref}$. The model is trained to reconstruct the full scene using $\mathcal{L}_{\text{flow}}$ (Eq. 5), conditioned on $Z_{t}$, $Z_{ref}$, and $Z_{c}$.
Spatial Style Encoder.
@@
Independently, we attach a new Style Branch ($Z_{style}$) to the same frozen, pre-trained FLUX.11-dev [19] base model and train only its LoRA adapters ($\Delta W_{\text{style}}$). This parallel process ensures the style representations are learned completely independently from the identity representations. The model is trained on a diverse corpus of $70,000$ images, including $40,000$ stylized scenes from OmniStyle [50], $15,000$ from StyleBooth [11], and $15,000$ real-world images. We train the model on a style-aware inpainting task. Given an image latent $Z_{i}$ and a binary mask $M$, we define the style context as the unmasked tokens $Z_{\text{style}}=Z_{i}\odot(1-M)$ and the noisy target tokens as $Z_{t}$ (the masked region $Z_{i}\odot M$, noised). The model is trained to denoise $Z_{t}$ conditioned on $Z_{\text{style}}$ and a text prompt $Z_{c}$.
4.3 Insert In Style: Style Aware Reference Composition
To formulate the complete compositional framework, we integrate the pre-trained, Identity Branch ($\Delta W_{\text{ref}}$) and Style Branch ($\Delta W_{\text{style}}$) into the FLUX.1-dev [19] base model. As these branches are optimized for disparate objectives, standard inference-time adapter mixing [45] yields sub-optimal performance. Consequently, we introduce a novel set of trainable LoRA adapters, $\Delta W_{\theta}$, applied to the main branch ($Z_{t}$). This compositional stage is trained using our proposed Insert In Style dataset, comprising 115115k curated samples (see Sec. 3).
However, naïvely fine-tuning the unified model inherently compromises the generalization capabilities of the individual constituent branches. This degradation arises fundamentally from the joint optimization process itself, forcing previously isolated semantic spaces to update simultaneously leads to severe feature entanglement, independent of the composite data distribution. To mitigate this structural interference, we propose Masked Attention to explicitly decouple the Identity and Style branches within the multi-modal attention mechanism. Furthermore, directly finetuning LoRA weights for complex compositional tasks inherently risks over-entangling the structural and stylistic conditions, leading to catastrophic forgetting of the base model’s broad generative priors. Therefore, we incorporate a prior preservation term into our training objective (Eq.11).
Structural Attention Mask